# Differentially Private Clustering in Data Streams

### v2

@@ 4 Bicriteria Approximation in Continual Release Setting
1:Privacy parameter $\varepsilon$, Stream ${\mathcal{S}}$ of points $x_{1},\ldots,x_{T}\in{\mathbb{R}}^{d}$
2:Initialize($\varepsilon$):
3:$\varepsilon^{\prime}:=\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$
4:Parallel quadtrees $Q_{1},\ldots,Q_{\log(k)}$ such that: each quadtree $Q_{q}$ has $\log(\Lambda)$ levels with the bottom level having grid size $\Theta(1)$
5:for $0\leq\ell\leq\log(\Lambda)$ and $1\leq q\leq\log(k)$ do
6:     Initialize $\mathbf{DPFindCenters}_{\ell,q}$ of $\mathbf{DPFindCenters}(\varepsilon^{\prime})$
7:Set of candidate centers ${\mathcal{F}}:=\emptyset$
8:Update($x_{t}$):
9:for $0\leq\ell\leq\log(\Lambda)$ and $1\leq q\leq\log(k)$ do
10:     ^$\hat{{\mathcal{F}}}_{t}\leftarrow\mathbf{Update}(x_{t})\text{ of }\mathbf{DPFindCenters}_{\ell,q}$$\triangleright$ See Algorithm 3
11:     ^${\mathcal{F}}\leftarrow{\mathcal{F}}\cup\hat{{\mathcal{F}}}_{t}$
12:Output ${\mathcal{F}}$
@@
1:Privacy parameter $\varepsilon^{\prime}$, Stream of points $x_{1},\ldots,x_{T}$
2:Set of candidate centers ^$\hat{{\mathcal{F}}}_{t}$ at every timestep $t\in[T]$
3:Initialize($\varepsilon^{\prime}$):
4:$\varepsilon^{\prime}\leftarrow\varepsilon^{\prime}$triangleright where $\varepsilon^{\prime}:=\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$
5:$w=O(k)$
[…]
7:$\hat{T}_{1}=0,\ldots,\hat{T}_{w}=0$$\triangleright$ DP Count for the size of hash bucket
8:Initialize $\textbf{BM}_{1},\ldots,\textbf{BM}_{w}$ of $\textbf{BinaryMechanism}(T,\varepsilon^{\prime})$$\triangleright$ See Theorem 5 [22]
[…]
10:Update($x_{t}$):
11:Initialize ^$\hat{{\mathcal{F}}}_{t}=\emptyset$
12:Let $x_{t}$ be mapped to cell $\mathbf{c}^{*}$triangleright Recall DPFindCenters is initialized per level $\ell$ of quadtree instance $q$
13:for $p=1,\ldots,L$, where $L:=\log(k^{2})$ run in parallel do
14:     for $j\in[w]$ do $\triangleright$ Update the DP count of each hash bucket
15:         if $j=h(\mathbf{c}^{*})$ then
16:              $\hat{T}_{j}\leftarrow\textbf{Update}(1)\text{ of }\textbf{BM}_{j}$
17:         else
18:              $\hat{T}_{j}\leftarrow\textbf{Update}(0)\text{ of }\textbf{BM}_{j}$
19:     for $j\in[w]$ do $\triangleright$ Update the DP HHs of each hash bucket
20:         if $h(\mathbf{c}^{*})=j$ then
[…]
22:         else
[…]
24:         for $\mathbf{c}\in H$ do
25:              if $\hat{f}(\mathbf{c})\geq\frac{\theta}{1000}\cdot\hat{T}_{h(\mathbf{c})}$ then
26:                  Add centerpoint of $\mathbf{c}$ to ^$\hat{{\mathcal{F}}}_{t}$ as a center
27:Return ^$\hat{{\mathcal{F}}}_{t}$
@@
Our bicriteria approximation algorithm is given by Algorithm 2 which initializes $\log(k)$ parallel instances of randomly shifted quadtrees. Each input point $x_{t}$ is assigned to a cell in every level of every quadtree. For a fixed quadtree $1\leq q\leq\log(k)$ and fixed level $0\leq\ell\leq\Lambda$, the subroutine DPFindCenters (see Algorithm 3) returns a candidate set of centers ^$\hat{{\mathcal{F}}}_{t}$ which is added to the current set of candidate set of centers ${\mathcal{F}}$.
[…]
4.1 Proof of Theorem 9
@@
There are $\log(\Lambda)$ levels per quadtree, thus point $x_{t}$ is a member of $\log(\Lambda)$ cells in total. As there are $2\log^{2}(k)$ parallel processes — considering $\log(k)$ quadtrees and $\log(k^{2})$ parallel processes (see Line 13 of Algorithm 3) per quadtree — a single point participates in $2\log\Lambda\log^{2}(k)$ total calls to DP-HH. Since DP-HH is DP under the continual release setting (by Theorem 6), assigning each DP-HH instance a privacy budget of $\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$ preserves $(\varepsilon/2)$-DP by sequential composition (Item 1 of Theorem 4).
@@
[17] Let ${\mathcal{B}}$ be an $\ell_{\infty}$ ball of radius $r$ contained in $[-\Lambda,\Lambda]^{d}$ (it forms a $d$-dimensional cube with each side length $2r$). Then for a randomly shifted quadtree and any level $\ell$ with grid size at least $r^{\prime}\geq 2r$, ${\mathcal{B}}$ is split by the grid in each dimension $j\in[d]$ independently with probability $\frac{2r}{r^{\prime}}$.
@@
=$\sum^{\log(\Lambda)}_{\ell=0}(\text{grid length at level }\ell)^{2}\cdot n_{G_{\ell}}$
@@
$=\sum^{T}_{t=1}(\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ uncovered by }G_{\ell})]+\mathbbm{1}[(x_{t}\text{ uncoveredby }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})])$
$=\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ uncovered by }G_{\ell})]+\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})]$
The first sum in the above expression can be upper bounded by $n_{G_{\ell}}$, thus it remains to bound the second sum. In order to bound the second sum, we will need some properties of good cells that are hashed to buckets in DPFindCenters, the proofs of these claims can be found in Appendix A. In the sequel, we denote $N_{\ell,\mathbf{c}}$ as the number of points in the cell $\mathbf{c}$ at level $\ell$ and the hash buckets as ${\mathcal{B}}_{j}$ where $j\in[w]$. For simplicity, we consider the number of hash buckets $w:=40k$. We first show that for any good cell $\mathbf{c}$, it is unlikely that the bucket it is hashed to contains another good cell $\mathbf{c}^{\prime}\neq\mathbf{c}$.
Claim 2.
@@
Observe that for a fixed hash bucket ${\mathcal{B}}_{j}$, any cell $\mathbf{c}$ such that $N_{\ell,\mathbf{c}}\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})$ qualifies as a $\theta$-heavy hitter since $N_{\ell,\mathbf{c}}\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})\geq\theta|{\mathcal{B}}_{j}|$ (by Claim 3). In particular, for good cell $\mathbf{c}_{y}$ such that $N_{\ell,\mathbf{c}_{y}}=y$, if $\mathbf{c}_{y}$ is an $\theta$-HH then $y\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})$ or $y\geq\frac{\theta n_{G_{\ell}}}{20k}$. We formalize this intuition in the claim below where we use the accuracy guarantees of DP-HH given by Theorem 6 to characterize the good cells that are reported as $\theta$-HHs.
Claim 4.
@@
$\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})]\leq\theta n_{G_{\ell}}+\frac{8k\log(\Lambda)\log^{2}(k)}{\varepsilon\gamma_{h}}\operatorname*{poly}\left(\log\left(\frac{T\cdot k\cdot 2^{\ell}}{\theta\gamma_{h}}\right)\right)$
Thus by combining Claim 5 with our observation about the first sum being upper bounded by $n_{G_{\ell}}$ in the decomposition of $\sum^{T}_{t=1}\mathbbm{1}[x_{t}\text{ uncovered by }H_{\ell,t}]$, we obtain our desired statement in Lemma 8. ∎
Note that we have shown Lemma 8 is true with probability at least $1-\frac{12}{k}$, for a fixed level. Since we have $\log(\Lambda)$ many levels in a specific quadtree, and $\log(k)$ many quadtree instances in parallel — we can boost our probability of success to be sufficiently high. It remains to bound the total $k$-means cost for the set of centers ${\mathcal{F}}$ output by our algorithm. Combining Equation 4.1, Equation 31 and Equation 4.1 along with Lemma 8, we obtain the following.
@@
Claim 6 (Size of ${\mathcal{F}}$ in Theorem 9).
For all $j\in[w]$, suppose the size of the hash bucket ${\mathcal{B}}_{j}$ is $|{\mathcal{B}}_{j}|=\Omega(\frac{\log(\Lambda)\log^{3}(k)}{\varepsilon}\log^{2.5}T)$ then with high probability, the total number of historical heavy hitters at the end of the stream is $O({k}\log^{2}(k)\log(\Lambda)\log T)$.
@@
The algorithm runs independent instances of the DP-HH algorithm for each bucket of each level in each instantiation of the quadtree, thus it is sufficient to first show that for a fixed quadtree $Q$, a fixed level $\ell$, and a fixed bucket ${\mathcal{B}}_{j}$ where $j\in[w]$, the total number of historical heavy hitters is at most $O(\frac{(1+\gamma_{h})}{{\theta}}\log T)$.
Let the timestamps of points that end up in ${\mathcal{B}}_{j}$ be $t_{i}=2^{i}$, where $0\leq i\leq\log(T)$. Let the state of the hash bucket at time step $t$ be ${\mathcal{B}}^{(t)}_{j}$. We set the failure probability in Theorem 5 as $\xi:=\frac{1}{k^{2}}$. From Theorem 5 we know that with probability $1-\frac{1}{k^{2}}$, the DP count of the hash bucket $\hat{T}_{j}$ at timestep $t$ has additive error $O(\frac{\log(\Lambda)\log^{3}(k)}{\varepsilon}\log^{2.5}(T))$. Thus for a fixed timestamp $t$, if $|{\mathcal{B}}^{(t)}_{j}|=\Omega(\frac{\log(\Lambda)\log^{3}(k)}{\varepsilon}\log^{2.5}(T))$, then we can see that a cell $\mathbf{c}$ is added to ${\mathcal{F}}_{t}$ only if $\hat{f}(\mathbf{c})>\frac{2\theta}{1000}|{\mathcal{B}}^{(t)}_{j}|$. Recall from Condition 1 of Theorem 6 that $\hat{N}_{\ell,\mathbf{c}}\in(1\pm\gamma_{h})N_{\ell,\mathbf{c}}$. Thus if $\mathbf{c}\in{\mathcal{F}}_{t}$ and $|{\mathcal{B}}^{(t)}_{j}|=\Omega(\frac{\log(\Lambda)\log^{3}(k)}{\varepsilon}\log^{2.5}(T))$ then it must be the case that with probability $1-\frac{1}{k^{2}}$, $N_{\ell,\mathbf{c}}\geq\frac{{2\theta}}{1000(1+\gamma_{h})}\|{\mathcal{B}}^{(i)}_{j}\|_{1}\geq\frac{{2\theta}}{1000(1+\gamma_{h})}{t_{i-1}}$.
Now, suppose for a contradiction, that the number of heavy hitters between $t_{i-1}$ and $t_{i}$ is at least $\frac{1000(1+\gamma_{h})}{{\theta}}$. Then for each such cell $\mathbf{c}$, we have that $N_{\ell,\mathbf{c}}\geq\frac{{2\theta}}{1000(1+\gamma_{h})}{t_{i-2}}$. Since there are at least $\frac{2000(1+\gamma_{h})}{{2\theta}}$ such cells, this implies that the total number of points between $t_{i-1}$ and $t_{i}$ is $\geq\frac{2\theta}{1000(1+\gamma_{h})}{t_{i-2}}\frac{2000(1+\gamma_{h})}{{2\theta}}=2^{i}=t_{i}$, which is a contradiction. Thus there must be at most $\frac{1000(1+\gamma_{h})}{{\theta}}$ cells marked as heavy hitters between consecutive intervals, and since there are $\log(T)$ such intervals, we have that the total number of historical $\ell_{1}$ heavy hitters for a fixed bucket is $O(\frac{(1+\gamma_{h})}{{\theta}}\log T)$.
Boosting the success probability over $O(k)$ buckets, $\log(k^{2})$ parallel processes, $\log(\Lambda)$ quadtree levels, and $\log(k)$ parallel processes of the quadtree instantiation, accounting for the additional number of historical HHs, and taking $\theta$ and $\gamma_{h}$ as appropriate constants, we obtain the claim as stated. ∎
5 DP Merge And Reduce Algorithm

### v3

@@ Appendix B Bicriteria Approximation in Continual Release Setting
Data: Privacy parameter $\varepsilon$, Stream $\mathcal{S}$} of points $x_{1},\ldots,x_{T}\in{\mathbb{R}}^{d}$
0.75em
Initialize($\varepsilon$):
$\varepsilon^{\prime}:=\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$;
Parallel quadtrees $Q_{1},\ldots,Q_{\log(k)}$ such that:
each quadtree $Q_{q}$ has $\log(\Lambda)$ levels with the bottom level having grid size $\Theta(1)$;
for $0\leq\ell\leq\log(\Lambda)$ and $1\leq q\leq\log(k)$ do
  Initialize $\mathbf{DPFindCenters}_{\ell,q}$ of $\mathbf{DPFindCenters}(\varepsilon^{\prime})$;
Set of candidate centers $\mathcal{F}$$\emptyset$;
0.75em Update($x_{t}$):
for $0\leq\ell\leq\log(\Lambda)$ and $1\leq q\leq\log(k)$ do
 $\hat{{\mathcal{F}}}_{t}\leftarrow\mathbf{Update}(x_{t})\text{ of }\mathbf{DPFindCenters}_{\ell,q}$;
 $\triangleright$ See algorithm 3
 ${\mathcal{F}}\leftarrow{\mathcal{F}}\cup\hat{{\mathcal{F}}}_{t}$;
Output $\mathcal{F}$};
@@
Data: Privacy parameter $\varepsilon^{\prime}$, Stream of points $x_{1},\ldots,x_{T}$, level of quadtree $\ell$
Initialize($\varepsilon^{\prime}$):
$\varepsilon^{\prime}\leftarrow\varepsilon^{\prime}$;
$\triangleright$ where $\varepsilon^{\prime}:=\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$
$w=O(k)$ ;
Hash function $h:[2^{\ell}]\to[w]$ s.t.  $\forall\text{ cells }\mathbf{c},\forall j\in[w],\Pr[h(\mathbf{c})=j]=\frac{1}{w}$;
$\hat{T}_{1}=0,\ldots,\hat{T}_{w}=0$;
$\triangleright$ DP Count for the size of hash bucket
Initialize $\textbf{BM}_{1},\ldots,\textbf{BM}_{w}$ of $\textbf{BinaryMechanism}(T,\varepsilon^{\prime})$;
$\triangleright$ See Theorem 9 [24]
Initialize $\textbf{DP-HH}_{1},\ldots,\textbf{DP-HH}_{w}$ of $\textbf{DP-HH}(T,\varepsilon^{\prime})$;
$\triangleright$ See Theorem 10 [27]
Update($x_{t}$):
Initialize $\hat{{\mathcal{F}}}_{t}=\emptyset$;
Let $x_{t}$ be mapped to cell $\mathbf{c}^{*}$;
$\triangleright$ DPFindCenters is initialized per level $\ell$ of quadtree instance $q$
for $p=1,\ldots,L$, where $L:=\log(k^{2})$ run in parallel do
 for $j\in[w]$ do
    $\triangleright$ Update the DP count of each hash bucket
    if $j=h(\mathbf{c}^{*})$ then
       $\hat{T}_{j}\leftarrow\textbf{Update}(1)\text{ of }\textbf{BM}_{j}$;
    else
       $\hat{T}_{j}\leftarrow\textbf{Update}(0)\text{ of }\textbf{BM}_{j}$
 for $j\in[w]$ do
    $\triangleright$ Update the DP $H$s of each hash bucket
    if $h(\mathbf{c}^{*})=j$ then
       $\hat{f},H\leftarrow$ Update($\mathbf{c}^{*}$) of $\textbf{DP-HH}_{j}$;
    else
       $\hat{f},H\leftarrow$ Update($\perp$) of $\textbf{DP-HH}_{j}$;
    for cell $\mathbf{c}\in H$ do
       if $\hat{f}(\mathbf{c})\geq\frac{\theta}{1000}\cdot\hat{T}_{h(\mathbf{c})}$ then
            Add centerpoint of $\mathbf{c}$ to $\hat{{\mathcal{F}}}_{t}$ as a center;
          $\triangleright$ $\theta$ is the $H$ threshold parameter and set to be an appropriate constant
Return $\hat{{\mathcal{F}}}_{t}$;
@@
Our bicriteria approximation algorithm is given by algorithm 2 which initializes $\log(k)$ parallel instances of randomly shifted quadtrees. Each input point $x_{t}$ is assigned to a cell in every level of every quadtree. For a fixed quadtree $1\leq q\leq\log(k)$ and fixed level $0\leq\ell\leq\Lambda$, the subroutine DPFindCenters (see algorithm 3) returns a candidate set of centers $\hat{{\mathcal{F}}}_{t}$ which is added to the current set of candidate set of centers $\mathcal{F}$}.
The DPFindCenters subroutine (see algorithm 3) finds the approximate heaviest $O(k)$ cells in a fixed level of a fixed quadtree. It achieves this by first hashing the cell containing the current point to a bucket, note that there are $w:=O(k)$ many buckets. For each hash bucket $j\in[w]$, the algorithm maintains a continual release $\theta$-heavy hitter instance $\textsf{DP-HH}_{j}$. We use the $\ell_{1}$-heavy hitter algorithm from [27] as DP-HH — it returns a set $H$ of $\theta$-heavy hitters and their approximate counts $\hat{f}(\mathbf{c})$ for all $\mathbf{c}\in H$. Since we are storing the centerpoints of all the cells marked as heavy hitters as candidate centers, we need to ensure that we do not store too many false positives, i.e., cells whose counts are much smaller than $\theta$ times the length of the hashed substream. To address this challenge, we have an additional pruning step that eliminates any cell $\mathbf{c}$ whose approximate count is less than $\Theta(\theta)\hat{T}_{h(\mathbf{c})}$ where $\hat{T}_{h(\mathbf{c})}$ denotes the DP count of each hash bucket $j\in[w]$ at timestep $t\in[T]$. We keep track of $\hat{T}_{j}$ via an instance of the Binary Mechanism [24] denoted as $\textbf{BM}_{j}$ for each $j\in[w]$. Finally, only the centerpoints of cells that pass this pruning step are added as candidate centers to the set $\hat{{\mathcal{F}}}_{t}$.
B.1 Proof of Theorem 4
@@
There are $\log(\Lambda)$ levels per quadtree, thus point $x_{t}$ is a member of $\log(\Lambda)$ cells in total. As there are $2\log^{2}(k)$ parallel processes — considering $\log(k)$ quadtrees and $\log(k^{2})$ parallel processes per quadtree — a single point participates in $2\log\Lambda\log^{2}(k)$ total calls to DP-HH. Since DP-HH is DP under the continual release setting (by Theorem 10), assigning each DP-HH instance a privacy budget of $\frac{\varepsilon}{4\log(\Lambda)\log^{2}(k)}$ preserves $(\varepsilon/2)$-DP by sequential composition (Item 1 of Theorem 8).
@@
[18] Let $\mathcal{B}$} be an $\ell_{\infty}$ ball of radius $r$ contained in $\Lambda,\Lambda]^{d}$ (it forms a $d$-dimensional cube with each side length $2r$). Then for a randomly shifted quadtree and any level $\ell$ with grid size at least $r^{\prime}\geq 2r$, $\mathcal{B}$} is split by the grid in each dimension $j\in[d]$ independently with probability $\frac{2r}{r^{\prime}}$.
@@
$\sum^{\log(\Lambda)}_{\ell=0}(\text{grid length at level }\ell)^{2}\cdot n_{G_{\ell}}$
@@
$=\sum^{T}_{t=1}(\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ uncovered by }G_{\ell})]$
$+\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})])$
$=\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ uncovered by }G_{\ell})]$
$+\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})]$
The first sum in the above expression can be upper bounded by $n_{G_{\ell}}$, thus it remains to bound the second sum. In order to bound the second sum, we will need some properties of good cells that are hashed to buckets in DPFindCenters, the proofs of these claims can be found in Appendix E. In the sequel, we denote $N_{\ell,\mathbf{c}}$ as the number of points in the cell $\mathbf{c}$ at level $\ell$ and the hash buckets as ${\mathcal{B}}_{j}$ where $j\in[w]$. For simplicity, we consider the number of hash buckets $w:=40k$. We first show that for any good cell $\mathbf{c}$, it is unlikely that the bucket it is hashed to contains another good cell $\mathbf{c}^{\prime}\neq\mathbf{c}$.
Proposition 2.
@@
Observe that for a fixed hash bucket ${\mathcal{B}}_{j}$, any cell $\mathbf{c}$ such that $N_{\ell,\mathbf{c}}\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})$ qualifies as a $\theta$-heavy hitter since $N_{\ell,\mathbf{c}}\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})\geq\theta|{\mathcal{B}}_{j}|$ (by Proposition 3). In particular, for good cell $\mathbf{c}_{y}$ such that $N_{\ell,\mathbf{c}_{y}}=y$, if $\mathbf{c}_{y}$ is an $\theta$-HH then $y\geq\theta\cdot 2(y+\frac{n_{G_{\ell}}}{40k})$ or $y\geq\frac{\theta n_{G_{\ell}}}{20k}$. We formalize this intuition in the claim below where we use the accuracy guarantees of DP-HH given by Theorem 10 to characterize the good cells that are reported as $\theta$-$H$s.
Proposition 4.
@@
$\sum^{T}_{t=1}\mathbbm{1}[(x_{t}\text{ uncovered by }H_{\ell,t})\land(x_{t}\text{ covered by }G_{\ell})]$
$\leq\theta n_{G_{\ell}}+\frac{8k\log(\Lambda)\log^{2}(k)}{\varepsilon\gamma_{h}}\operatorname*{poly}\left(\log\left(\frac{T\cdot k\cdot 2^{\ell}}{\theta\gamma_{h}}\right)\right)$
Thus by combining Proposition 5 with our observation about the first sum being upper bounded by $n_{G_{\ell}}$ in the decomposition of $\sum^{T}_{t=1}\mathbbm{1}[x_{t}\text{ uncovered by }H_{\ell,t}]$, we obtain our desired statement in Lemma 5. ∎
Note that we have shown Lemma 5 is true with probability at least $1-\frac{12}{k}$, for a fixed level. Since we have $\log(\Lambda)$ many levels in a specific quadtree, and $\log(k)$ many quadtree instances in parallel — we can boost our probability of success to be sufficiently high. It remains to bound the total $k$-means cost for the set of centers $\mathcal{F}$} output by our algorithm. Combining Section B.1, eq. 4 and Section B.1 along with Lemma 5, we obtain the following.
@@
Lemma 7 (Size of $\mathcal{F}$} in Theorem 4).
For all $j\in[w]$, suppose the size of the hash bucket ${\mathcal{B}}_{j}$ is $|{\mathcal{B}}_{j}|=\Omega(\frac{\log(\Lambda)\log^{3}(k)}{\varepsilon}\log^{2.5}T)$ then with high probability, the total number of heavy hitters at the end of the stream is $O({k}\log^{2}(k)\log(\Lambda)\log T)$.
@@
The algorithm runs independent instances of the DP-HH algorithm for each bucket of each level in each instantiation of the quadtree, thus it is sufficient to first show that for a fixed quadtree $Q$, a fixed level $\ell$, and a fixed bucket ${\mathcal{B}}_{j}$ where $j\in[w]$, the total number of heavy hitters is at most $O(\frac{(1+\gamma_{h})}{{\theta}}\log T)$.
[…]
Now, suppose for a contradiction, that the number of heavy hitters between $t_{i-1}$ and $t_{i}$ is at least $\frac{1000(1+\gamma_{h})}{{\theta}}$. Then for each such cell $\mathbf{c}$, we have that $N_{\ell,\mathbf{c}}\geq\frac{{2\theta}}{1000(1+\gamma_{h})}{t_{i-2}}$. Since there are at least $\frac{2000(1+\gamma_{h})}{{2\theta}}$ such cells, this implies that the total number of points between $t_{i-1}$ and $t_{i}$ is $\geq\frac{2\theta}{1000(1+\gamma_{h})}{t_{i-2}}\frac{2000(1+\gamma_{h})}{{2\theta}}=2^{i}=t_{i}$, which is a contradiction. Thus there must be at most $\frac{1000(1+\gamma_{h})}{{\theta}}$ cells marked as heavy hitters between consecutive intervals, and since there are $\log(T)$ such intervals, we have that the total number of $\ell_{1}$ heavy hitters for a fixed bucket is $O(\frac{(1+\gamma_{h})}{{\theta}}\log T)$.
Boosting the success probability over $O(k)$ buckets, $\log(k^{2})$ parallel processes, $\log(\Lambda)$ quadtree levels, and $\log(k)$ parallel processes of the quadtree instantiation, accounting for the additional number of heavy hitters, and taking $\theta$ and $\gamma_{h}$ as appropriate constants, we obtain the claim as stated. ∎
Appendix C DP Merge And Reduce Algorithm